OpenStudy (anonymous):

Solve the system of equations using matrices. Use Gaussian elimination with back-substitution. 3x + 5y - 2w = -13 2x + 7z - w = -1 4y + 3z + 3w = 1 -x + 2y + 4z = -5

OpenStudy (anonymous):

these are my answers

OpenStudy (arindameducationusc):

4 variables, 4 equations..... you can solve each value...

OpenStudy (anonymous):

not really but i can try

OpenStudy (arindameducationusc):

@Vocaloid would you like to try this one?

OpenStudy (anonymous):

3x + 5y - 2w = -13 x=3

OpenStudy (anonymous):

2x + 7z - w = -1 x=2

OpenStudy (anonymous):

sorry x= -2

OpenStudy (anonymous):

am i getting these right

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

-x + 2y + 4z = -5 x=4

OpenStudy (arindameducationusc):

no not like that.... i will solve it, wait....

OpenStudy (anonymous):

ohh ok

OpenStudy (arindameducationusc):

you have to solve one equation for one varialble, here2x=-13-5y+2w now put this in second equation.... 2x+7z-w=-1 =>-13-5y+2w+7z-w=-1

OpenStudy (arindameducationusc):

You have to do like this okay?

OpenStudy (anonymous):

yeah im lost

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

is the answer a

OpenStudy (arindameducationusc):

suppose x+y=15 and 2x+8y=10 we solve x and y. same way we are solving 4 variables